OpenStudy (anonymous):

Evaluate the function y = x+5/x^2+7x+12 at x=-5 x=1. Write undefined if appropriate.

OpenStudy (saifoo.khan):

\[y = \frac{x+5}{x^2 + 7x + 12} \] @ x= -5, \[y = \frac{(-5)+5}{(-5)^2 + 7(-5) + 12} \] ---------------------------------------- @ x = 1 \[y = \frac{(1)+5}{(1)^2 + 7(1) + 12} \]

OpenStudy (anonymous):

undefined at -5... and @1 it is 3/10